OpenStudy (anonymous):

Kindly confirm please.. Is tan^2 (3x-1) tan3x can be reduced to tan^3 (9x^2 - 3x) ??

OpenStudy (anonymous):

is tan^2 (3x-1) tan3x =tan^2 (3x-1) tan(3x) ? or =tan^2 (3x-1) tan^3(x) ?

OpenStudy (anonymous):

the first one. :D

OpenStudy (anonymous):

then it's reduced down as much as it can be, the only way you could do the tan^3 is if all the angles were the same, and if they were all the same, then it would look like: tan^3(3x-1) make sense?